Solve the IPV using Operators: y+ y= x,y(0)=1, y(0) -0
Answer #1

tå y+y=x, 4101=1,4%) = 0 & {y+y} = 4(x) 52 715) +5 Y/5) = — Y15) = Sestris5315+) 5 + 55-sti Thin taring Lt, we have ylul -